OpenStudy (anonymous):

A cash register contains $341 dollars in $20, $5, and $2 bills. There are 28 bills in all, and three more twos than fives. How many bills of each kind are there? x= number of $20 x+y+z=28 bills y= number of $5 20x+5y+2z=341 z=number of $2 ? (Can you tell me how to get this? I have it, but don't know how to get it)

OpenStudy (anonymous):

There are 28 bills in all. Which means x+y+z=28, right? Then you know that the register contains $341, 1x = $20, 1y = $5 and 1z = $2 so, making an equation out of that gives you 20x+5y+2z=341

OpenStudy (anonymous):

What about the equation for three more twos than fives?

OpenStudy (anonymous):

I know it y=z -3, but how do you get this equation?

OpenStudy (anonymous):

There are three more twos than fives. There are x twenties y fives y+3 twos (Twos are THREE MORE than fives.) or x twenties z twos z-3 fives (Fives are THREE LESS than twos.)

OpenStudy (anonymous):

Thank you so much!!!